The sequences are defined using recursion formulas. Write the first four terms of each sequence. a1 = 12 and an = an-1 + 4 for n ≧ 2.

Answer by Edwin McCravy(20055)

a1 = 12 and an = an-1 + 4 for n ≧ 2.

----------------------------------------------------------

a1 = 12

a2 = a2-1 + 4 = a1 + 4 = 12 + 4 = 16

a3 = a3-1 + 4 = a2 + 4 = 16 + 4 = 20

a4 = a4-1 + 4 = a3 + 4 = 20 + 4 = 24


Edwin